OpenStudy (rose4825):

How many points are there on a taxicab circle with radius 4? 22 16 18 20 My answer: 16

OpenStudy (rose4825):

@jim_thompson5910 @mathstudent55

jimthompson5910 (jim_thompson5910):

in `taxicab geometry`, you can only move about on the integer coordinates. You cannot move in the space between the integer coordinate points because those spaces represent houses where the taxi can't drive anyways, a `taxicab circle` is where all points are equally distant from a fixed point. In this example, the red points are all equally distant from the blue point https://upload.wikimedia.org/wikipedia/commons/thumb/d/de/TaxicabGeometryCircle.svg/214px-TaxicabGeometryCircle.svg.png it turns out that there are 16 points shown and the radius is 4. So you are correct

jimthompson5910 (jim_thompson5910):

when I say "distance" I mean the distance the car drives you cannot move diagonally

jimthompson5910 (jim_thompson5910):

In that latest screenshot, I show 2 points being 4 spaces away from the blue point

OpenStudy (rose4825):

Alright, this one was confusing me but I think I got the answer. Complete question with drawing is here: http://prntscr.com/bchtsy Rectangle ABCD is a golden rectangle. What is AB if BC is 10? 6.2 –16.2 5.4 4.8 My answer: 6.2

jimthompson5910 (jim_thompson5910):

We can't have a negative length, so cross off choice B

jimthompson5910 (jim_thompson5910):

BC is 10 AB is some length smaller than 10 (based on the answer choices I see)

jimthompson5910 (jim_thompson5910):

The golden ratio is approximately 1.61803399

jimthompson5910 (jim_thompson5910):

the idea is that we divide the length and height to get 1.61803399

jimthompson5910 (jim_thompson5910):

BC/AB = 1.61803399 10/AB = 1.61803399 10 = 1.61803399*AB 1.61803399*AB = 10 AB = 10/1.61803399 AB = 6.18033988272398 which is approximately 6.2

jimthompson5910 (jim_thompson5910):

so you got it correct

OpenStudy (rose4825):

Wow I did not think I got that one right, but thanks. Okay, last one: http://prntscr.com/bchv49 I have no idea on this one because these graphs are really confusing me

OpenStudy (rose4825):

Maybe 1 and 2, but I do not know

jimthompson5910 (jim_thompson5910):

This is a tough one. A `Hamiltonian Circuit` is basically where you visit each node on a closed path. You start at one node, visit the rest, then come back to the starting node. You visit each node EXACTLY ONE time Anyways, you just have to do this through trial and error. Here is what I came up with

OpenStudy (rose4825):

So you're saying choices 2 and 3?

jimthompson5910 (jim_thompson5910):

`An Euler path is a path that uses every edge of a graph exactly once` Source: https://www.math.ku.edu/~jmartin/courses/math105-F11/Lectures/chapter5-part2.pdf so in choice 2, we don't use path AC or CA in choice 3, my screenshot isn't using path AC or CA so choices 2 and 3 aren't euler paths

OpenStudy (rose4825):

Oh okay, I get your point. I guessed on this one because I had no idea what to pick, but I think based on your explanation I got another question wrong. I check that one in a second.

OpenStudy (rose4825):

http://prntscr.com/bchxzn I chose Hamiltonian circuits for this one, but I do not think this is correct anymore :\

OpenStudy (rose4825):

It seems like there are no Hamiltonian circuits or Euler paths

jimthompson5910 (jim_thompson5910):

I agree with your answer

OpenStudy (rose4825):

So I was correct with my initial answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (rose4825):

Thank you so much!! You are a wonderful helper.

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (rose4825):

The questions with the choices was incorrect for future references, it was actually choices 3 and 4 :)